Nova Companys total overhead cost at various levels of activity are presented below: Total Machine-Overhead Month April May June July Hours Cost 55,000 205,600 45,000179,600 65,000 75,000 ş 231,600 ş 257,600 Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 45,000 machine-hour level of activity is Utilities (variable) Supervisory salaries (fixed) Maintenance (mixed) Total overhead cost $67,500 49,000 63,100 $179,600 Nova Companys management wants to break down the maintenance cost into its variable and fixed cost elements Required 1. Estimate how much of the $257,600 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $257600 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.) 2. Using the high-low method, estimate a cost formula for maintenance in the form Y=a + bX. 3. Express the companys total overhead cost in the form Y=a + bX. 4. What total overhead cost would you expect to be incurred at an activity level of 50,000 machine-hours?
Answer #1
1
Maintenance cost in July = 257600-(67500/45000*75000)-49000= 96100
2
Machine hours Maintenance cost
High level 75000 96100
Low level 45000 63100
Change 30000 33000
Variable cost = 33000/30000= $1.1
Fixed cost = 96100-(75000*1.1) = $13600
Y = 13600+1.1X
3
Variable cost Fixed cost
Utilities 1.5
Supervisory salaries 49000
Maintenance cost 1.1 13600
2.6 62600
Y = 62600+2.6X
4
Total overhead cost = 62600+(50000*2.6)= 192600